OpenStudy (anonymous):

Solve by completing the square: x^2+2x-7=0 Need some help on understanding this please.

OpenStudy (luigi0210):

Do you know how to start?

OpenStudy (anonymous):

I thought you got it. D: What part do you still need help with? :) http://openstudy.com/study#/updates/527c028ae4b0de4fd4770081

OpenStudy (idku):

add 8 to both sides to make the left side into a perfect square.

OpenStudy (solomonzelman):

Yep add 8 to both sides\[x^2+2x-7=0\]\[x^2+2x-7+8=0+8⇒ x^2+2x+1=8\] is 'x^2+2x+1' a perfect square?

OpenStudy (solomonzelman):